OpenStudy (mackenzie2013):

After 6 months the simple interest earned annually on an investment of $6,000 was $463. Find the interest rate to the nearest tenth of a percent.

OpenStudy (anonymous):

I = Prt with T used in years... I = interest p = principal r = interest rate t = time

OpenStudy (mackenzie2013):

it would be 0.5

OpenStudy (mackenzie2013):

@JuanitaM

OpenStudy (jdoe0001):

hmmmm

OpenStudy (anonymous):

I got a slightly different answer

OpenStudy (mackenzie2013):

Dont you have to convert the 6 months into a year?

OpenStudy (anonymous):

You were provided only I, P , and T remember T = 6/12 = 1/2

OpenStudy (anonymous):

Yes you do that is .5 or 1/2

OpenStudy (anonymous):

The question wants you to find R the rate..so plug everything in and tell me what you got for R

OpenStudy (jdoe0001):

\(\qquad \textit{Simple Interest Earned}\\ I = Prt\qquad \begin{cases} I=\textit{interest earned}\to &\$463\\ P=\textit{original amount deposited}\to& \$6000\\ r=rate\\ t=years\to &6 \end{cases} \\ \quad \\ \textit{solve for "r"}\)

OpenStudy (mackenzie2013):

So i take 6,000= 463 x R X 0.5

OpenStudy (anonymous):

aah...P = principal which is 6000 i - interest which is 463

OpenStudy (anonymous):

Annual investment is termed they are using for the principal amount

OpenStudy (mackenzie2013):

I'm pretty confused

OpenStudy (anonymous):

You were on the right path, you just mixed up principal and interest amount

OpenStudy (anonymous):

Simple interest is I = PRT

OpenStudy (anonymous):

what amount did they tell you was interest? principal? and you got T correct at .5

OpenStudy (mackenzie2013):

6,000 and 463?

OpenStudy (anonymous):

yes, now plug it all in solve for R like you did the first time...you were close don't get frustrated...

OpenStudy (anonymous):

so 463 = 6000 * r *.5

OpenStudy (mackenzie2013):

15%?

OpenStudy (anonymous):

I actually got .16 and the question said to round to the nearest tenth of a percent..

OpenStudy (mackenzie2013):

15.4%

OpenStudy (mackenzie2013):

Thank you for your help!!

OpenStudy (anonymous):

Yes,my calculator was on 2 decimals...I will take your 15.4%..Glad I could be of assistance! Happy Studying